# Quantum-enhanced sensing of spin-orbit coupling without fine-tuning

### v1

@@ 4.3 Thermal probe
Figure 3: Thermal probe. Probe state is the thermal state of the Hamiltonian in Eq. (11) with $t{=}1$, $\theta{=}0$, $\phi{=}0$, $B{=}0.01t$, and $\alpha_{y}{=}\alpha_{z}$. (a) QFI vs. temperature in the non-interacting case for different system sizes. As thermal energy surpasses the gap, QFI falls of universally as ${\sim}1/T$. (b) Variation of QFI with both $U$ and $T$ for $L{=}6$.
So far we have discussed the cases where the probe is taken to be the ground state which is the equilibrium state at zero temperature. In practice however, the sensors generally operate at finite temperature and hence, it is necessary to analyze the effect of temperature on the sensing capability. When the temperature is smaller than the energy gap i.e. $k_{B}T{<}\Delta$ ($k_{B}$ is the Boltzmann constant), we expect the ground state description to hold. Higher temperatures result in a mixed state with increasing contributions from higher energy states. The QFI then starts to decrease and typically collapses universally with an algebraic decay with temperature [74]. In Fig. 3(a), we show this behavior in the non-interacting case for two system sizes $L=100,200$. The ground state QFI value is retained up to a temperature ${\sim}\Delta/k_{B}$, after which the QFI is found to fall off as $T^{-1}$. In the repulsive interacting case, we observe that the energy gap decreases with $U$, and hence the temperature at which the QFI starts to fall decreases with $U$ (see Fig. 3(b)).
@@ 5 Multi-parameter sensing
Figure 4: Multi-parameter sensing. Probe state is the single particle ground state of the Hamiltonian in Eq. (6) with $t{=}1$, $\theta{=}0$, $\phi{=}0$, and $B{=}0.01t$. (a) The $(\alpha_{z},\alpha_{z})$ component of the QFIM $\mathcal{F}^{Q}$ across a range of $\alpha_{z}$ and $\alpha_{z}$ for $L{=}100$. Four points chosen in the parameter space for scaling analysis are $(\alpha_{y},\alpha_{z})=(0.05,0.1),(0.9,0.05),(0.9,0.9),(0.05,0.9)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling of $(\mathcal{F}^{Q})_{\alpha_{z},\alpha_{z}}$ at the four points. (c) The $(\alpha_{y},\alpha_{y})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{y}}$ at the four points. (e) The $(\alpha_{y},\alpha_{z})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{z}}$ at the four points.
In this section, we consider a more general scenario where we do not presume prior knowledge of the SOC parameters i.e. the Rashba fields in any direction. The goal is to estimate both $\alpha_{y}$ and $\alpha_{z}$ simultaneously, making the problem multi-parametric. To analyze the system’s sensitivity to changes in these parameters, we compute the multi-parameter QFIM $\mathcal{F}^{Q}$, which provides a comprehensive description of the ultimate precision limits for estimating multiple parameters concurrently. In Fig. 4(a) we shows the component $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ as a function of the two parameters. Figure 4(b) shows that $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ reaches super-linear scaling in a wide range of $\alpha_{z}$ and $\alpha_{y}$ and can also attain close to Heisenberg scaling. Similar observation also holds for the component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{y}}$, as shown in Figs. 4(c) and (d). However, the sensitivity also depends on the off-diagonal component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ that captures the correlation between the estimates of $\alpha_{z}$ and $\alpha_{y}$. Figures 4(e) and (f) show that the scaling of $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ also stays close to quadratic across the whole range. From this analysis we can also infer that the quantum enhancement of the sensing of $\alpha_{y}$ when $\alpha_{z}$ is known, can reach the Heisenberg limit as well.
@@ 7 Conclusion
YB acknowledges support from National Natural Science Foundation of China (Grant No. 12404551) and the China Postdoctoral Science Foundation (Grant No. 2024M750339). SS acknowledges support from National Natural Science Foundation of China (Grant No. W2433012). AB acknowledges support from National Natural Science Foundation of China (Grants Nos. 12050410253, 92065115, and 12274059) and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L).

### v2

@@ 4.3 Thermal probe
Figure 3: Thermal probe. Probe state is the thermal state of the Hamiltonian in Eq. (10) with $t{=}1$, $B{=}0.01t$, $\alpha_{y}{=}\alpha_{z}$. (a) QFI vs. temperature in the non-interacting case for different system sizes. As thermal energy surpasses the gap, QFI falls of universally as ${\sim}1/T$. (b) Variation of QFI with both $U$ and $T$ for $L{=}6$.
So far we have discussed the cases where the probe is taken to be the ground state which is the equilibrium state at zero temperature. In practice however, the sensors generally operate at finite temperature and hence, it is necessary to analyze the effect of temperature on the sensing capability. When the temperature is smaller than the energy gap i.e. $k_{B}T{<}\Delta$ ($k_{B}$ is the Boltzmann constant), we expect the ground state description to hold. Higher temperatures result in a mixed state with increasing contributions from higher energy states. The QFI then starts to decrease and typically collapses universally with an algebraic decay with temperature [74]. In Fig. 3(a), we show this behavior in the non-interacting case for two system sizes $L=100,200$. The ground state QFI value is retained up to a temperature $T{\sim}\Delta/k_{B}$, after which the QFI is found to fall off as $T^{-1}$. In the repulsive interacting case, we observe that the energy gap decreases with $U$, and hence the temperature at which the QFI starts to fall decreases with $U$ (see Fig. 3(b)).
@@ 5 Multi-parameter sensing
Figure 4: Multi-parameter sensing. Probe state is the single particle ground state of the Hamiltonian in Eq. (6) with $t{=}1$, and $B{=}0.01t$. (a) The $(\alpha_{z},\alpha_{z})$ component of the QFIM $\mathcal{F}^{Q}$ across a range of $\alpha_{z}$ and $\alpha_{z}$ for $L{=}100$. Four points chosen in the parameter space for scaling analysis are $(\alpha_{y},\alpha_{z})=(0.05,0.1),(0.9,0.05),(0.9,0.9),(0.05,0.9)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling of $(\mathcal{F}^{Q})_{\alpha_{z},\alpha_{z}}$ at the four points. (c) The $(\alpha_{y},\alpha_{y})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{y}}$ at the four points. (e) The $(\alpha_{y},\alpha_{z})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{z}}$ at the four points.
In this section, we consider a more general scenario where we do not presume prior knowledge of the SOC parameters i.e. the Rashba fields in any direction. The goal is to estimate both $\alpha_{y}$ and $\alpha_{z}$ simultaneously, making the problem multi-parametric. To analyze the system’s sensitivity to changes in these parameters, we compute the multi-parameter QFIM $\mathcal{F}^{Q}$, which provides a comprehensive description of the ultimate precision limits for estimating multiple parameters concurrently. In Fig. 4(a) we shows the component $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ as a function of the two parameters. Fig 4(b) shows that $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ reaches super-linear scaling in a wide range of $\alpha_{z}$ and $\alpha_{y}$ and can also attain close to Heisenberg scaling. Similar observation also holds for the component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{y}}$, as shown in Figs. 4(c) and (d). From this analysis we can also infer that the quantum enhancement of the sensing of $\alpha_{y}$ when $\alpha_{z}$ is known, can reach the Heisenberg limit as well. In general, however, the sensitivity depends on the off-diagonal component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ that captures the correlation between the estimates of $\alpha_{z}$ and $\alpha_{y}$. Fig 4(e) and (f) show that the scaling of $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ also stays close to quadratic across the whole range.
@@ 7 Conclusion
YB acknowledges support from National Natural Science Foundation of China (Grant No. 12404551) and the China Postdoctoral Science Foundation (Grant No. 2024M750339). AB acknowledges support from National Natural Science Foundation of China (Grants Nos. 12050410253, 92065115, and 12274059) and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). SS acknowledges support from National Natural Science Foundation of China (Grant No. W2433012).